# Exploration v.s. Exploitation: Rethinking RLVR through Clipping, Entropy, and Spurious Reward

### v1

@@ Conclusion.
Combining Eq.˜18 and Eq.˜21, we obtain the desired result
${\mathbb{E}}[{\mathcal{H}}(\pi_{c})-{\mathcal{H}}(\pi_{u})]$ (22)
$\leq$ $-\tfrac{1-2^{1-G}}{2G}\Phi(\pi_{\textnormal{old}})\eta^{2}+{\mathbb{E}}[R(\eta)]+c(p)G\left(\rho\delta_{\rm eff}-\tfrac{X_{\max}}{2}(G-1)p\right)$ (23)
$=$ $-\tfrac{1-2^{1-G}}{2G}\Phi(\pi_{\textnormal{old}})\eta^{2}+{\mathbb{E}}[R(\eta)]-G\pi_{\min}\left(\log(pe^{\eta/(2\pi_{\min})})\right)_{-}\left(\rho\delta_{\rm eff}-\tfrac{X_{\max}}{2}(G-1)p\right).$ (24)
@@
In this remark, we will show that under practical settings, the assumption, i.e., Eq.˜17, in Theorem˜4.3 is indeed satisfied and the extra term ensure the entropy is decreasing in expectation. Recall the parameters we used or observed in experiments: $G=16$, $\eta=5\times 10^{-7}$, $|\mathcal{V}|\approx 150000$, $\pi_{\min}=10^{-7}$, ${\epsilon}=0.2$, $\rho\approx 0.001$, $\delta\approx 10$, and we have freedom to choose $p=2\pi_{\min}=2\times 10^{-7}$. By Lemma˜C.5, $\mathbb{P}(\mathcal{B})\leq 0.0071$. We can choose $\hat{\pi}=10^{-6}$, then $|\mathbb{E}[R_{c}(\eta)]|\leq 4.32\times 10^{-9}$. The third term in LABEL:eq:HcHu-bound-final is $6.72\times 10^{-8}$. Thus, we require $\Phi(\pi_{\rm old})>-8.05\times 10^{6}$ to ensure $\mathbb{E}[\Delta\mathcal{H}]<0$. However, we compute that $\Phi_{\min}=-2.23\times 10^{-6}$, thus the condition $\Phi(\pi_{\rm old})>-8.05\times 10^{6}$ always hold and we can conclude $\mathbb{E}[\Delta\mathcal{H}]<0$.
Lemma C.5.
Let $\mathcal{B}:=\left\{\exists i:\pi(\mathbf{y}^{(i)})=\pi_{\min},\,A_{i}>0,\,r_{u}(\mathbf{y}^{(i)})\leq 1+{\epsilon}\right\}$ and
$Z_{\max}(\eta):=1+\tfrac{\eta}{2}\exp\left(\tfrac{\eta}{2\pi_{\min}}\right),\quad A_{0}^{*}:=\tfrac{G\pi_{\min}}{\eta}\log\left((1+{\epsilon})Z_{\max}(\eta)\right).$
Let $\mathcal{V}_{\min}:=\{a\in\mathcal{V}:\pi(a)=\pi_{\min}\}$ and $\mu_{\min}:=\pi(\mathcal{V}_{\min})\leq|\mathcal{V}|\pi_{\min}$. Define the collision event among $\pi_{\min}$-tokens:
$\operatorname{Coll}_{\min}:=\{\exists a\in\mathcal{V}_{\min}:N_{a}\geq 2\},\quad N_{a}:=\sum_{i=1}^{G}\mathbf{1}\{Y^{(i)}=a\}.$
Then
$\mathbb{P}(\mathcal{B})\leq G\mu_{\min}\mathbb{P}(0<A_{1}\leq A_{0}^{*})+\tbinom{G}{2}|\mathcal{V}|\pi_{\min}^{2}.$
Proof.
Recall that $|\tilde{A}(a)|\leq 1/(2\pi_{\min})$. Write
$Z(\eta)-1=\sum_{a\in\mathcal{V}}\pi(a)(e^{\eta\tilde{A}(a)}-1)=\sum_{a:\tilde{A}(a)>0}\pi(a)(e^{\eta\tilde{A}(a)}-1),$
since $e^{x}-1\leq 0$ for $x\leq 0$. For $x\geq 0,e^{x}-1\leq xe^{x}$. Hence
$Z(\eta)-1\leq\sum_{a:\tilde{A}(a)>0}\pi(a)\eta\tilde{A}(a)e^{\eta\tilde{A}(a)}.$
Using $e^{\eta\tilde{A}(a)}\leq\exp(\eta/(2\pi_{\text{min}}))$, we get
$Z(\eta)-1\leq\eta\exp\left(\tfrac{\eta}{2\pi_{\min}}\right)\sum_{a:\tilde{A}(a)>0}\pi(a)\tilde{A}(a)\leq\tfrac{1}{2}\exp\left(\tfrac{\eta}{2\pi_{\min}}\right).$
Thus, $Z(\eta)\leq Z_{\max}(\eta)$. Decompose $\mathcal{B}\subset\left(\mathcal{B}\cap\operatorname{Coll}_{\min}^{c}\right)\cup\operatorname{Coll}_{\min}$. On $\operatorname{Coll}_{\min}^{c}$, if $\pi(\mathbf{y}^{(i)})=\pi_{\min}$, then the token $\mathbf{y}^{(i)}$ appears exactly once in the batch, hence $\tilde{A}(\mathbf{y}^{(i)})=\tfrac{A_{i}}{G\pi_{\mathrm{min}}}$. If additionally $r_{u}(\mathbf{y}^{(i)})\leq 1+{\epsilon}$, then by definition
$\tfrac{\exp(\eta\tilde{A}(\mathbf{y}^{(i)}))}{Z(\eta)}\leq 1+{\epsilon}\implies\exp(\eta\tilde{A}(\mathbf{y}^{(i)}))\leq(1+{\epsilon})Z(\eta)\leq(1+{\epsilon})Z_{\max}(\eta),$
Taking logs and substituting $\tilde{A}(Y^{(i)})=A_{i}/(G\pi_{\text{min }})$ gives
$A_{i}\leq\tfrac{G\pi_{\min}}{\eta}\log((1+{\epsilon})Z_{\max}(\eta))=A_{0}^{*}.$
Therefore,
$\mathcal{B}\cap\operatorname{Coll}_{\min}^{c}\subset\{\exists i:\pi(\mathbf{y}^{(i)})=\pi_{\min},0<A_{i}\leq A_{0}^{*}\}.$
By union bound,
$\mathbb{P}(\mathcal{B}\cap\text{Coll}_{\min}^{c})\leq\sum_{i=1}^{G}\mathbb{P}\left(\pi(\mathbf{y}^{(i)})=\pi_{\min},0<A_{i}\leq A_{0}^{*}\right).$
Since $\mathbf{y}^{(i)}$ is independent of $A_{i}$,
$\mathbb{P}\left(\pi(\mathbf{y}^{(i)})=\pi_{\min},0<A_{i}\leq A_{0}^{*}\right)=\underbrace{\mathbb{P}\left(\pi(\mathbf{y}^{(i)})=\pi_{\min}\right)}_{=\mu_{\min}}\cdot\mathbb{P}\left(0<A_{1}\leq A_{0}^{*}\right),$
so $\mathbb{P}(\mathcal{B}\cap\operatorname{Coll}_{\min}^{c})\leq G\mu_{\min}\mathbb{P}(0<A_{1}\leq A_{0}^{*})$. For any fixed $a\in\mathcal{V}_{\min},\,N_{a}\sim\operatorname{Binomial}(G,\pi_{\min})$ and $\mathbb{P}\left(N_{a}\geq 2\right)\leq\tbinom{G}{2}\pi_{\min}^{2}$. Union bound over all $a\in\mathcal{V}_{\text{min }}$ yields
$\mathbb{P}(\mathrm{Coll}_{\min})\leq|\mathcal{V}_{\min}|\tbinom{G}{2}\pi_{\min}^{2}\leq V\tbinom{G}{2}\pi_{\min}^{2}.$
This completes the proof. ∎
@@ C.4 Missing Proofs in §5
Taking the conditional expectation of both sides of Eq.˜25 yields
@@
This together with Eq.˜26 yields the desired result. ∎

### v2

@@ C.3 Missing Proofs in §4
Combining Eq.˜29 and Eq.˜32, we obtain the desired result
$\begin{array}[]{cl}&{\mathbb{E}}[{\mathcal{H}}(\pi_{c})-{\mathcal{H}}(\pi_{u})]\\ \leq&-\tfrac{1-2^{1-G}}{2G}\Phi(\pi_{\textnormal{old}})\eta^{2}+{\mathbb{E}}[R(\eta)]+c(p)G\left(\rho\delta_{\rm eff}-\tfrac{X_{\max}}{2}(G-1)p\right)\\ =&-\tfrac{1-2^{1-G}}{2G}\Phi(\pi_{\textnormal{old}})\eta^{2}+{\mathbb{E}}[R(\eta)]-G\pi_{\min}\left(\log(pe^{\eta/(2\pi_{\min})})\right)_{-}\left(\rho\delta_{\rm eff}-\tfrac{X_{\max}}{2}(G-1)p\right).\end{array}$ (33)
@@
Lemma C.6.
Let $\mathcal{V}_{\leq\hat{\pi}}:=\{a\in\mathcal{V}:\pi_{\rm old}(a)\leq\hat{\pi}\}$ with some $\hat{\pi}\in[\pi_{\min},1)$. Define $\operatorname{Coll}_{\leq\hat{\pi}}:=\{\exists a\in\mathcal{V}_{\leq\hat{\pi}}:N_{a}\geq 2\}$ where $N_{a}:=\sum_{i=1}^{G}\mathbf{1}\{\mathbf{y}^{(i)}=a\}$. Then $\mathbb{P}(\operatorname{Coll}_{\leq\hat{\pi}})\leq\tbinom{G}{2}|\mathcal{V}|\hat{\pi}^{2}$. Furthermore, if $\hat{\pi}\leq\tfrac{G}{2(\sqrt{G}-1/\sqrt{G})}\pi_{\min}$, then on $\operatorname{Coll}_{\leq\hat{\pi}}^{c}$, we have $|\tilde{A}(a)|\leq\tfrac{1}{2\hat{\pi}}$ for all $a$.
Proof.
[…]
$\mathbb{P}(\operatorname{Coll}_{\leq\hat{\pi}})\leq\tbinom{G}{2}\sum_{a:\pi(a)\leq\hat{\pi}}\pi(a)^{2}\leq\tbinom{G}{2}\hat{\pi}\sum_{a:\pi(a)\leq\hat{\pi}}\pi(a)=\tbinom{G}{2}\hat{\pi}\mu_{\leq\hat{\pi}}$,
and using $\mu_{\leq\hat{\pi}}\leq|\mathcal{V}|\hat{\pi}$ gives the final bound $\mathbb{P}(\operatorname{Coll}_{\leq\hat{\pi}})\leq\tbinom{G}{2}|\mathcal{V}|\hat{\pi}^{2}$. On $\operatorname{Coll}_{\leq\hat{\pi}}^{c}$, if $\pi_{\rm old}(a)\leq\hat{\pi}$, then the token $a$ appears at most once in the batch, hence $\tilde{A}(a)=\tfrac{A_{i}}{G\pi(a)}$ or $\tilde{A}(a)=0$. For the former one, by Lemma˜2.4, we have $|\tilde{A}(a)|\leq\tfrac{\sqrt{G}-1/\sqrt{G}}{G\pi(a)}\leq\tfrac{\sqrt{G}-1/\sqrt{G}}{G\pi_{\min}}$. By assumption $\hat{\pi}\leq\tfrac{G}{2(\sqrt{G}-1/\sqrt{G})}\pi_{\min}$, one concludes that $|\tilde{A}(a)|\leq\tfrac{1}{2\hat{\pi}}$. ∎
@@
In this remark, we will show that under practical settings, the assumption, i.e., Eq.˜28, in Theorem˜4.3 is indeed satisfied and the extra term ensure the entropy is decreasing in expectation. Recall the parameters we used or observed in experiments: $G=16$, $\eta=5\times 10^{-7}$, $|\mathcal{V}|\approx 150000$, $\pi_{\min}=10^{-7}$, ${\epsilon}=0.2$, $\rho\approx 0.001$, $\delta\approx 10$, and we choose the threshold $p=\hat{\pi}=2\pi_{\min}=2\times 10^{-7}$. By Theorem˜4.3, we compute $X_{\max}=10.98$, $M(p)=2.29$, $\delta_{\rm eff}=9.74$, and $c(p)=-1.29\times 10^{-6}$. Thus, the clipping correction term (the third term) in the upper bound of $\mathbb{E}[\Delta\mathcal{H}]$ is $2.01\times 10^{-7}$. Furthermore, $c_{G}\eta^{2}=7.81\times 10^{-15}$ and $\Phi(\pi)\geq\Phi_{\min}=-2.23\times 10^{6}$. Finally, define $B:=\operatorname{Coll}_{\leq\hat{\pi}}$ in Lemma˜C.6, applying Corollary˜C.2, we have
$|\mathbb{E}[R(\eta)]|\leq(1-q_{\hat{\pi}})C(\hat{\pi})+q_{\hat{\pi}}C(\pi_{\min}),$
where $q_{\hat{\pi}}\leq 7.2\times 10^{-7}$ by Lemma˜C.6, $C(\hat{\pi})=3.43\times 10^{-8}$, and $C(\pi_{\min})=5.31\times 10^{-7}$. Hence, $|\mathbb{E}[R(\eta)]|\leq 3.43\times 10^{-8}$. Combining all numbers, we have
$\mathbb{E}[\Delta\mathcal{H}]\leq(-2.23\times 10^{6})\times(-7.81\times 10^{-15})+3.43\times 10^{-8}-2.01\times 10^{-7}<-1.49\times 10^{-7}<0.$
Thus, we can conclude $\mathbb{E}[\Delta\mathcal{H}]<0$.
@@ C.4 Missing Proofs in §5
Taking the conditional expectation of both sides of Eq.˜34 yields
@@
This together with Eq.˜35 yields the desired result. ∎